OpenStudy (anonymous):

x2 - 49 = 0 (factoring) Need help not answers

OpenStudy (anonymous):

Rule is x^2-a^2=(x-a)(x+a) so you can use this rule...

OpenStudy (anonymous):

400^2m use the formula A=s^2 explain how would it be used to find the length od side of lot?

OpenStudy (anonymous):

Use: \[\large a^2 - b^2 = (a+b)(a-b)\] Here a = x and b = 7..

OpenStudy (anonymous):

Just take the square root of 400 you will get the length...

OpenStudy (anonymous):

See, \(Area = Side^2\) So, \(Side = \sqrt{Area}\) \(Side = \sqrt{400}\) What is the square root of 400??

mathslover (mathslover):

I just liked your spirit @jezzieljademacaraeg ...keep it up but please reply to the response given by waterineyes to ur question .. wht is the square root of 400 ?

OpenStudy (anonymous):

Thanks for helping @mathslover

mathslover (mathslover):

Pleasure of an amby ...